# A Data-Aware Fourier Neural Operator for Modeling Spatiotemporal Electromagnetic Fields

### v1

@@ 3.2 Training
The proposed DA-FNO model eliminates the CFL condition, enabling larger time-step iterations while simultaneously generating all components of the electromagnetic field at each iteration. Before training, the model’s temporal resolution is set to 15 times the FDTD time step ($\Delta t_{model}=15\Delta t_{FDTD}$, corresponding to $m=15$ in Fig. 1). We sample the full-field data $\left(E_{z},H_{x},H_{y}\right)$ from datasets at every 15th FDTD time step and scale the magnetic field components by the intrinsic impedance of free space, $Z_{0}=\sqrt{\mu_{0}/\epsilon_{0}}=120\pi\ \Omega$, to match the magnitude of the electric field. Subsequently, each data sample is normalized to the range $\left[-1,1\right]$, motivating the use of a tanhtanh activation function in the model’s output layer. These preprocessing steps promote numerical stability, balanced feature scaling, and improved training generalization.
@@
Due to the energy convergence termination criterion ($\delta=10^{-4}$) in the model simulation, the simulation duration—and consequently the temporal length of the generated data—varies across different geometries. To facilitate batch processing and accelerate training, each sample within a mini-batch is zero-padded to a uniform length $T$, and a mask is defined to record the padded entries for subsequent loss computation. For each mini-batch, the initial five time slices serve as the input $\mathbf{x}$, while the remaining T5T-5 time slices constitute the label data $\mathbf{y}$. The model iteratively predicts the field components for T5T-5 steps, and all predicted components are concatenated to form $\mathbf{\hat{y}}$. The loss function for backpropagation is then computed using $\mathbf{y}$, $\mathbf{\hat{y}}$, and the $\mathbf{mask}$:
$\mathcal{J}(\mathbf{\hat{y},\mathbf{y},\mathbf{mask}})=\sum_{s,t}(\mathbf{mask}\odot(\frac{1}{3}\sum_{c=1}^{3}\mathcal{L}_{2}(\mathbf{\hat{y}},\mathbf{y}))),$ (5)
@@
where $s$,$t$ and $c$ denotes the sample index, time index and component index, respectively. The $\mathbf{mask}$ is a 2D tensor of 0s and 1s, where 1 indicates valid data and 0 denotes padding, excluding the influence of padded regions during backpropagation. $\mathcal{L}_{2}(\mathbf{\hat{y}},\mathbf{y})$ represents the relative $L_{2}$ loss:
[…]
@@
The loss function $\mathcal{J}$ computes the relative $L_{2}$ error across all field components and time steps, thereby preventing the model from discounting later time steps with lower magnitudes and encouraging more consistent prediction accuracy throughout the entire temporal sequence.
All training was conducted on NVIDIA A6000 and A100 GPUs.
3.3 Results
To evaluate the performance on training and testing datasets, we use the relative $L_{1}$ error:
$\mathcal{L}_{1}\left({\hat{y}},{y}\right)=\frac{{\left\|{\hat{y}}-{y}\right\|_{1}}}{{\left\|{y}\right\|_{1}}}.$ (7)
throughout.
[…]
We first evaluate the training performance of the DA-FNO models under different threshold values $\theta$, and compare them with that of the Conv-FNO and vanilla FNO models. The Conv-FNO model refers to the vanilla FNO with the convolutional processing module from DA-FNO, while retaining the original rectangular mode selection strategy. Notably, the number of selected modes in Conv-FNO and vanilla FNO is slightly higher than that of DA-FNO with $\theta=90\%$.
Through several training runs, we observe that initial conditions have a significant impact on the training dynamics and final performance of the models. Given this sensitivity, we conduct repeated training trials with random weight initializations to ensure that the evaluation results are representative and not biased by outlier cases. For each model, these trials are performed using the first 500 samples of the full dataset, which are partitioned into a training subset (samples 1-450) and a test subset (samples 451-500). Among all trials, three runs with stable and consistent loss descent curves are selected for comparison. All models are trained for 1000 epochs under identical hyperparameter settings.
The results are presented in Fig. 6. For DA-FNO models, increasing the threshold $\theta$ leads to a consistent downward trend in both training and test errors, indicating that retaining more spectral modes improves the model’s ability to simulate electromagnetic fields. However, when $\theta$ exceeds 0.9, the remaining modes (as shown in the spectral distribution of Fig. 4(b)) exhibit very low amplitudes, and even a slight increase in $\theta$ leads to a substantial rise in the number of selected modes. Therefore, we do not consider higher values of $\theta$.
Compared with the Conv-FNO model, the DA-FNO ($\theta=90\%$) model achieves a slightly lower minimum error in the Fig. 6, suggesting that the data-aware mode selection strategy is more effective than the rectangular strategy in capturing long-term electromagnetic dynamics. Additionally, both DA-FNO and Conv-FNO models significantly outperform the vanilla FNO models, especially in terms of test error, demonstrating that the incorporation of convolutional processing greatly enhances both model accuracy and generalization.
Figure 7: (a) Comparison between partial outputs ($H_{y}$, from $t=5\Delta t$ to $22\Delta t$) of the DA-FNO model and FDTD for a single sample. (b) The last two time-step results in (a) with individually tailored colormaps. (c) Comparative scatter plot of the normalized energy for all test samples of the DA-FNO model and FDTD. The black dashed line represents $y=x$.
Based on the above experiment, the DA-FNO model with $\theta=90\%$ is employed and trained on the entire dataset, with $\delta=10^{-4}$ as the convergence criterion. The final results show a average relative $L_{1}$ error of 8.43% on the training set and 9.82% on the testing set. A partial output (from $t=5\Delta t$ to $t=22\Delta t$) of one sample with a relative $L_{1}$ error of 9.86% is illustrated in Fig. 7(a). A TM-polarized plane wave interacts with the geometry upon incidence. As the wave propagates out of the simulation domain, the energy within the region gradually decays. It can be observed from the figure that the results obtained by FDTD and the DA-FNO model are highly similar. For clearer visual comparison, the last two time-step results shown in Fig. 7(a) are re-visualized in Fig. 7(b) using individually tailored colormaps. Despite the presence of numerous high-frequency variations in the figure, the outputs of FDTD and the model exhibit strong similarity. Moreover, we compare the normalized energy at each time step of the outputs from the model and FDTD across all test samples, as shown in the Fig. 7(c). Each point represents a specific time step of a test sample. Comparison with the reference line $y=x$ reveals that the model-predicted energies closely match those computed by the FDTD in the low-energy regime. In the intermediate range, the model tends to overestimate the energy, while in the high-energy regime, it slightly underestimates it, indicating a larger variation in the model’s predictions. Despite the variation, the predicted values remain well-aligned with the reference, demonstrating an overall acceptable level of consistency between the two approaches. This consistency ensures that the model’s predictions are sufficiently reliable to allow the simulation to terminate at the correct time steps based on energy monitoring.
Figure 8: Variation of the Ez value at the central point of the simulation domain predicted by the DA-FNO model and FDTD under different geometries. The blue lines represent the FDTD predictions, and the orange dots represent the DA-FNO model predictions.
In addition, Fig. 8 presents the variations of the $E_{z}$ field at the central point of the simulation domain for 12 random samples, with the geometry of each sample displayed in the bottom-right corner of each subplot. Across various randomized geometries and simulation durations, the predictions of the DA-FNO model (orange scatter points) align closely with that of the FDTD (blue curves), indicating high accuracy in capturing temporal waveforms. Similarly, since the temporal waveform at each point is available, it is also feasible to derive corresponding frequency-domain information or energy distributions. The accuracy of such calculations depends on the temporal resolution $\Delta t_{model}$ of the simulation. A coarse time step may lead to inaccuracies in the computed frequency-domain representations, while a finer time step yields more precise results. Overall, given access to time-domain data, it is possible to obtain frequency-domain characteristics through appropriate post-processing techniques.
[…]
Figure 9: Average relative $L_{1}$ error (%) at different central wavelengths and convergence criteria.
Table 1: Average Relative $L_{1}$ Error (%) versus Central Wavelengths ($\lambda_{0}/\text{nm}$) and Convergence Criteria ($\delta$)
$\delta\backslash\lambda_{0}$ 1530 1535 1540 1545 1550 1555 1560 1565 1570
$10^{-2}$ 7.77 7.64 7.55 7.50 7.50 7.54 7.63 7.77 7.84
$10^{-3}$ 8.65 8.51 8.41 8.34 8.33 8.36 8.44 8.57 8.73
$10^{-4}$ 10.17 10.02 9.91 9.84 9.82 9.83 9.90 10.01 10.16
The DA-FNO model is capable of overcoming the CFL constraint, enabling faster simulations. To evaluate simulation speed, we test the DA-FNO model with $\theta=90\%$, the in-house FDTD, and the Meep-based FDTD [25] with different numbers of samples. All experiments are conducted on an Intel(R) Xeon(R) Gold 6346 CPU (16 cores, 32 threads) and an NVIDIA GeForce RTX 4090 GPU. For conventional FDTD methods, GPU-based parallel simulation of individual samples is inefficient when the computational domain is small, as data transfer overhead between the CPU and GPU can dominate the overall runtime. Therefore, in these experiments, the FDTD approaches are implemented with CPU-based multithreading instead, with the optimal number of threads determined empirically. In contrast, the neural network-based model naturally supports batch processing due to its architectural design, which makes it well-suited for execution on GPUs. We set the batch size to 200 in the model simulation and use O2 optimization for the C++-based in-house FDTD simulation. The Meep-based FDTD simulations are executed with 16 threads to enable parallel processing.
As shown in Table 2, the simulation speed of the DA-FNO model increases with the number of samples when the sample size is below the batch size of 200, and stabilizes at approximately 0.034 s/sample when the sample size exceeds 200. This indicates that the DA-FNO model exhibits greater computational efficiency in large-sample scenarios. In contrast, the simulation speeds of the in-house FDTD and Meep FDTD remain nearly constant at approximately 0.3 s/sample and 0.254 s/sample, respectively, regardless of the sample size. Overall, when the sample size exceeds 200, the DA-FNO model achieves a speedup of approximately $7.5\times$ compared to the multithreaded Meep FDTD simulation.
Table 2: Simulation Time (s): DA-FNO vs. FDTD across Sample Sizes
Sample Size DA-FNO in-house FDTD Meep FDTD
50 2.20 14.94 12.78
100 3.87 31.70 25.60
200 6.81 61.37 50.56
300 10.28 96.19 75.77
4 Conclusion
In this work, we propose a novel Fourier neural operator tailored for photonic simulation, termed the Data-Aware Fourier Neural Operator (DA-FNO). The DA-FNO modifies the vanilla Fourier layer by incorporating convolutional operations to capture the interactions among different components of the electromagnetic field. Furthermore, it introduces a data-aware mode selection strategy within the Fourier layer, enabling adaptive selection of Fourier modes based on data characteristics, which enhances the capability for long-term electromagnetic simulations. By embedding the DA-FNO within an auto-regressive framework, we develop a novel neural-operator-based model capable of simultaneously predicting the temporal evolution of electromagnetic fields and generalizing across complex and randomized geometries. Moreover, it demonstrates strong wavelength generalization across the optical communication C-band. The DA-FNO model iteratively outputs all field components while monitoring energy variations to automatically terminate the simulation upon convergence. It achieves a speedup of $7.5\times$ compared to the conventional FDTD method, with an accuracy of 92.5%. Given its ability to predict the temporal evolution of electromagnetic fields, generalize across randomized geometries and the optical C-band, future work will focus on applying this FDTD surrogate model to the inverse design of photonic devices with time-dependent characteristics, such as those exhibiting dispersion. The model’s computational efficiency is particularly advantageous in the early stages of inverse design, where a large number of candidate configurations must be explored and high precision is not yet critical. We believe that the DA-FNO model offers a novel approach for photonic simulations and holds great potential as a powerful tool for the inverse design of devices with time-dependent characteristics.
Acknowledgment
This work is supported by High-performance Computing Platform of Peking University.

### v2

@@ Appendix B Network training
The proposed DA-FNO model eliminates the CFL condition, enabling larger time-step iterations while simultaneously generating all components of the electromagnetic field at each iteration. Before training, the model’s temporal resolution is set to $m$ times the FDTD time step ($\Delta t_{model}=m\Delta t_{FDTD}$). We sample the full-field data $\left(E_{z},H_{x},H_{y}\right)$ from datasets at every $m$th FDTD time step and scale the magnetic field components by the intrinsic impedance of free space, $Z_{0}=\sqrt{\mu_{0}/\epsilon_{0}}=120\pi\ \Omega$, to match the magnitude of the electric field. Subsequently, each data sample is normalized to the range $\left[-1,1\right]$, motivating the use of a tanhtanh activation function in the model’s output layer. These preprocessing steps promote numerical stability, balanced feature scaling, and improved training generalization.
@@
Due to the energy convergence termination criterion in the model simulation, the simulation duration—and consequently the temporal length of the generated data—varies across different geometries. To facilitate batch processing and accelerate training, each sample within a mini-batch is zero-padded to a uniform length $T$, and a mask is defined to record the padded entries for subsequent loss computation. For each mini-batch, the initial five time slices serve as the input $\mathbf{x}$, while the remaining T5T-5 time slices constitute the label data $\mathbf{y}$. The model iteratively predicts the field components for T5T-5 steps, and all predicted components are concatenated to form $\mathbf{\hat{y}}$. The loss function for backpropagation is then computed using $\mathbf{y}$, $\mathbf{\hat{y}}$, and the $\mathbf{mask}$:
$\mathcal{J}(\mathbf{\hat{y},\mathbf{y},\mathbf{mask}})=\sum_{n,t}(\mathbf{mask}\odot(\frac{1}{3}\sum_{c=1}^{3}\mathcal{L}_{2}(\mathbf{\hat{y}},\mathbf{y}))),$ (8)
@@
where $n$,$t$ and $c$ denotes the sample index, time index and component index, respectively. The $\mathbf{mask}$ is a 2D tensor of 0s and 1s, where 1 indicates valid data and 0 denotes padding, excluding the influence of padded regions during backpropagation. $\mathcal{L}_{2}(\mathbf{\hat{y}},\mathbf{y})$ represents the relative $L_{2}$ loss, defined as:
${\mathcal{L}_{2}}_{c,t}^{(n)}=\frac{\left(\sum_{i,j}\big|\hat{y}_{c,t}^{(n)}(i,j)-y_{c,t}^{(n)}(i,j)\big|^{2}\right)^{1/2}}{\left(\sum_{i,j}\big|y_{c,t}^{(n)}(i,j)\big|^{2}\right)^{1/2}}.$ (9)
@@
The loss function $\mathcal{J}$ computes the relative $L_{2}$ error across all field components and time steps, thereby preventing the model from discounting later time steps with lower magnitudes and encouraging more consistent prediction accuracy throughout the entire time-domain sequence.
All training was conducted on NVIDIA A6000 and A100 GPUs. More comprehensive description of the training process is provided in Supplementary material, Section 3.
Data Availability
The code developed for this study, together with the first 200 samples of the training dataset, is openly available at DA-FNO dataset (HuggingFace) . Additional datasets supporting the findings of this work are available from the corresponding author upon reasonable request.
Supplementary Material
Supplementary material associated with this article is provided as a separate file.